a)     Three year bonds are issued at face value of Sh100, 000 on Jan. 1, 2007, and a stated interest rate of 8%. Calculate the issue price of the bonds assuming a market interest rate of 6 %.( 5 marks)


1
Expert's answer
2022-03-24T16:14:17-0400

Interest earned per year= "\frac{8}{100} \times 100000 = sh.8000"

Present Value = "present\: value\: factor\times100000"

present value factor = "\frac{1}{(1+\frac{6}{100})^3}"

= 0.8396

PV= "0.8396\times100000=sh. 83960"

Present Value of interest Rates =  0.839619

Total Interest = "0.839619\times8000= 6,717"

Issue Price = "83960+6717"

=sh . 90677